HETAV PATEL

Grade 11,

8. The sum of all two digit odd numbers is (a) 2575 (b) 2475 (c) 2524 (d) 2425

8. The sum of all two digit odd numbers is
(a) 2575
(b) 2475
(c) 2524
(d) 2425

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: b Explaination:Reason: All two digit odd numbers are 11,13,15,… 99, which are in A.P. Since there are 90 two digit numbers of which 45 numbers are odd and 45 numbers are even ∴ Sum = 452[11 + 99] = 452 × 110 = 45 × 55 = 2475